Grade 11Mechanics

the distance traversed by a particle moving along a straight line is given by x=1801+50t2. the velocity at the end of 4s second

rajesh
8 Years agoGrade 11

1 Answer

Anshul Sharma
8 Years ago
X=180t+50t^2Therefore,integrate it you will get the velocity.V=180+100tNow, velocity at end of 4s isV=180+100(4)V=580m/s